Limited Resources
Maria Pajet, a regional sales representative for UniTec Systems, Inc., has been working about 80 hours per week calling on a total of 123 regular customers each month. Because of family and health considerations, she has decided to reduce her hours to a maximum of 160 per month. Unfortunately, this cutback will require Maria to turn away some of her regular customers or, at least, serve them less frequently than once a month. Maria has developed the following information to assist her in determining how to best allocate time:

Customer Classification
Large
Business
Small
Business
Individual
Number of customers 8 35 80
Average monthly sales per customer $2,500 $1,500 $600
Commission percentage 5% 8% 10%
Hours per customer per monthly visit 5.0 3.0 2.5


a. Develop a monthly plan that indicates the number of customers Maria should call on in each classification to maximize her monthly sales commissions.
b. Determine the monthly commissions Maria will earn if she implements this plan.

In the first table, calculate the commission per customer and per hour for each customer classification.

In the second table, complete the monthly call/visit plan that would maximize Maria's monthly sales commissions. Then determine Maria's monthly commissions if she implements this plan.

Commission
per customer
Commission
per hour
Large business Answer?? Answer??
Small business Answer?? Answer??
Individual Answer?? Answer??
Monthly Call/Visit Plan

Total number

of customers

Total number

of visit hours

Total

commissions

Large business Answer?? Answer?? Answer??
Small business Answer?? Answer?? Answer??
Individual Answer?? Answer?? Answer??
Grand totals: Answer?? Answer?? Answer??
Answer #1

Solution :- Maria paiet working about 80 hrs per week Calling Customers each month. On a total of 123 yegulay Large businessSmall business Commission per customer = 1,500*87. = 120 Commission per houx = 180/3.0 40 Individual Commission per customerIf 35 small business Costomers are visited then left out hours are :) 160-(35*3) .. » 55 Now, we need to maximize the 55 YestMonthly Cau/visit plan To tal no-of | Total no. of I Total Customers visit hours Cos tomers H0 2200 Large business 105 7560objective tonction .MAX No. of Costomers from each category X1 = 8 X2 <= 35 79<80 where . Xi - NO.of large business Costomers